The Harding Company manufactures skates. The company's Income statement for 20x1 is as follows: 


HARDING COMPANY Income Statement For the Year Ended December 31, 20x1 

Sales (11,700 skates @ $84 each)$982,800
Variable costs (11,700 skates at $37)432,900
Fixed costs320,000
Earnings before interest and taxes (EBIT)$229,900
Interest expense68,500
Earnings before taxes (EBT)$161,400
Income tax expense (30%)48,420
Earnings after taxes (EAT)$112,980


a. Compute the degree of operating leverage. 

b. Compute the degree of financial leverage

c. Compute the degree of combined leverage.

d. Compute the break-even point in units (number of skates). The Harding Company manufactures skates. The companys Income statement for 20x1 is as follows: $ HARDING COMPANY Income Statd. Compute the break-even point in units (number of skates). (Round your answer to the nearest whole number.) Break-even poin
Answer #1

Answer of Part a:

Degree of Operating Leverage = (Sales – Variable Cost) / (Sales – Variable Cost – Fixed Cost)
Degree of Operating Leverage = ($982,800 - $432,900) / ($982,800 - $432,900 - $320,000)
Degree of Operating Leverage = $549,900 / $229,900
Degree of Operating Leverage = 2.39

Answer of Part b:

Degree of Finance Leverage = EBIT / (EBIT – Interest)
Degree of Financial Leverage = $229,900 / ($229,900 - $68,500)
Degree of Financial Leverage = $229,900 / $161,400
Degree of Financial Leverage = 1.42

Answer of Part c:

Degree of Combined Leverage = Degree of Operating Leverage * Degree of Financial Leverage
Degree of Combined Leverage = 2.39 * 1.42
Degree of Combined Leverage = 3.40

Answer of Part d:

Contribution Margin per unit = Sales per unit – Variable cost per unit
Contribution Margin per unit = $84 - $37
Contribution Margin per unit = $47

Break Even Point in Units = Fixed Cost / Contribution Margin per unit
Break Even Point in Units = $320,000 / $47
Break Even Point in Units = 6,809 skates
Answer #2

SOLUTION :


a.


Degree of operating leverage (DOL)

= (Sales - Variable costs) / EBIT

= (982800 - 432900) / 229900

= 549900 / 229900

= 2.3919 

= 2.39  (ANSWER)


b.


Degree of financial leverage (DFL)

= EBIT / EBT 

= 229900 / 161400

= 1.4244

= 1.42 (ANSWER).


c.


Degree of combined leverage 

= Contribution margin / EBT  

= (982800 - 432900) / 161400

= 3.4071

= 3.41 (ANSWER).


d.


Break-even Point  in units = Fixed costs / Unit contribution margin

= 320000 / (84  - 37)

= 6808.51

= 6809  skates (ANSWER).


answered by: Tulsiram Garg